# Weakly Supervised Co-training with Swapping Assignments for Semantic Segmentation

### v1

@@ Abstract
Class activation maps (CAMs) are commonly employed in weakly supervised semantic segmentation (WSSS) to produce pseudo-labels. Due to incomplete or excessive class activation, existing studies often resort to offline CAM refinement, introducing additional stages or proposing offline modules. This can cause optimization difficulties for single-stage methods and limit generalizability. In this study, we aim to reduce the observed CAM inconsistency and error to mitigate reliance on refinement processes. We propose an end-to-end WSSS model incorporating guided CAMs, wherein our segmentation model is trained while concurrently optimizing CAMs online. Our method, Co-training with Swapping Assignments (CoSA), leverages a dual-stream framework, where one sub-network learns from the swapped assignments generated by the other. We introduce three techniques: i) soft perplexity-based regularization to penalize uncertain regions; ii) a threshold-searching approach to dynamically revise the confidence threshold; and iii) contrastive separation to address the coexistence problem. CoSA demonstrates exceptional performance, achieving mIoU of 76.2% and 51.0% on VOC and COCO validation datasets, respectively, surpassing existing baselines by a substantial margin. Notably, CoSA is the first single-stage approach to outperform all existing multi-stage methods including those with additional supervision.
@@ 1 Introduction
The objective of weakly supervised semantic segmentation (WSSS) is to train a segmentation model without relying on pixel-level labels but utilizing only weak and cost-effective annotations, such as image-level classification labels [3, 26, 50], object points [4, 45], and bounding boxes [16, 28, 57, 36]. In particular, image-level classification labels have commonly been employed as weak labels due to their minimal or negligible annotation effort involved [2, 60]. With the absence of precise localization information, image-level WSSS often necessitates the use of the coarse localization offered by class activation maps (CAMs) [72]. CAMs pertain to the intermediate outputs derived from a classification network. They can visually illustrate the activation regions corresponding to each individual class. Thus, they are often used to generate pseudo masks for training. However, CAMs suffer from i) Inconsistent Activation: CAMs demonstrate variability and lack robustness in accommodating geometric transformations of input images [60], resulting in inconsistent activation regions for the same input. ii) Inaccurate Activation: activation region accuracy is often compromised, resulting in incomplete or excessive class activation, only covering the discriminative object regions [1]. Despite enhanced localization mechanisms in the variants GradCAM [55] and GradCAM++ [7], they still struggle to generate satisfactory pseudo-labels for WSSS [60]. Thus, many WSSS works are dedicated to studying CAM refinement or post-processing [1, 15, 31].
@@
In general, WSSS methods [2, 65, 18, 50] comprise three stages: CAM generation, refinement, and segmentation training with pseudo-labels. This multi-stage framework is known to be time-consuming and complex as several models must be trained at different stages. In contrast, single-stage models [3, 70, 52], which include a unified network of all stages, are more efficient. They are trained to optimize the segmentation and classification tasks at the same time; however, their CAMs are not explicitly trained. As a result, they need refinement to produce high-quality pseudo-labels, often leveraging hand-crafted modules, such as CRF in [70], PAMR in [3], PAR in [52, 53]. As the refinement modules are predefined and offline, they decouple the CAMs from the primary optimization. When the refined CAMs are employed as learning objectives for segmentation, the optimization of the segmentation branch may deviate from that of the classification branch. Hence, it is difficult for a single-stage model to optimize its segmentation task while yielding satisfactory CAM pseudo-labels. This optimization difficulty underlies the inferior performance in single-stage approaches compared to multi-stage ones. Further, such hand-crafted refinement modules require heuristic tuning and empirical changes, thereby limiting their adaptability to novel datasets. Despite the potential benefits of post-refinement in addressing the aforementioned two issues associated with CAMs, which have been extensively discussed in WSSS studies, there has been limited exploration in explicit online optimization for CAMs.
@@
The absence of fully optimized CAMs is an important factor in the existing indispensability of this refinement. In this paper, we take a different approach, proposing a model that optimizes CAMs in an end-to-end fashion, resulting in reliable, consistent and accurate CAMs for WSSS without the necessity for subsequent refinements in two respects. First, we note that even though CAM is differentiable, it is not robust to variation. As the intermediate output of a classification model, CAMs are not fully optimized for segmentation purpose since the primary objective is to minimize classification error. This implies that within an optimized network, numerous weight combinations exist that can yield accurate classification outcomes, while generating CAMs of varying qualities. To investigate this, we conduct oracle experiments, training a classification model while simultaneously guiding the CAMs with segmentation ground truth. A noticeable enhancement in quality is observed in guided compared to vanilla CAMs, without compromising classification accuracy. Second, we demonstrate the feasibility of substituting the oracle with segmentation pseudo-labels (SPL) in the context of weak supervision. Consequently, we harness the potential of SPL for WSSS by co-training both CAMs and segmentation through mutual learning.
@@
We explore an effective way to substitute the CAM refinement process, i.e. guiding CAMs in an end-to-end fashion. Our method optimizes the CAMs and segmentation prediction simultaneously thanks to the differentiability of CAMs. To achieve this, we adopt a dual-stream framework that includes an online network (ON) and an assignment network (AN), inspired by self-supervised frameworks [5, 22]. The AN is responsible for producing CAM pseudo-labels (CPL) and segmentation pseudo-labels (SPL) to train the ON. Since CPL and SPL are swapped for supervising segmentation and CAMs, respectively, our method is named Co-training with Swapping Assignments (CoSA).
Our end-to-end framework enables us to leverage the quantified reliability of pseudo-labels for training online, as opposed to relying on offline hard pseudo-labels as the existing methodologies [2, 65, 15, 50]. Thus, our model can incorporate soft regularization driven by uncertainty, where the CPL perplexity is continually assessed throughout training. This regularization is designed to adaptively weight the segmentation loss, considering the varying reliability levels of different regions. By incorporating the soft CPL, CoSA enables dynamic learning at different time-steps rather than the performance being limited by the predefined CPL.
The threshold is a key parameter for generating the CPL [60, 50, 53]. It not only requires tuning but also necessitates dynamic adjustment to align with the model’s learning state at various time-steps. CoSA integrates threshold searching to dynamically adapt its learning state, as opposed to the commonly used hard threshold scheme [18, 13, 52]. With the proposed dynamic thresholding, we eliminate the laborious task of manual parameter tuning and enhance performance.
We further address a common issue in WSSS with CAMs, known as the coexistence problem, whereby certain class activations display extensive false positive regions that inaccurately merge the objects with their surroundings. In response, we introduce a technique to leverage low-level CAMs enriched with object-specific details to contrastively separate the foreground regions from the background.
The proposed CoSA greatly surpasses existing WSSS methods. Our approach achieves 76.2% and 75.1% mIoU on the VOC val and test splits, and 51.0% on the COCO val split, which are +5.1%, +3.9%, and +8.7% ahead of the previous single-stage SOTA [53].
@@
The contributions of this paper are as follows: 1) We are the first to propose SPL as a substitute for guiding CAMs in WSSS. We present compelling evidence showcasing its potential to produce more reliable, consistent and accurate CAMs. 2) We introduce a dual-stream framework with swapped assignments in WSSS, which co-optimizes the CAMs and segmentation predictions in an end-to-end fashion. 3) We develop a reliability-based adaptive weighting in accommodating the learning dynamics. 4) We incorporate threshold searching to automatically adjust the threshold, ensuring alignment with the learning state at different training time-steps. 5) We address the CAM coexistence issue and propose a contrastive separation approach to regularize CAMs. This greatly alleviates the coexistence problem, significantly enhancing the results of affected classes. 6) We demonstrate CoSA’s SOTA results on key challenging WSSS benchmarks, significantly surpassing existing methods. Our source code and model weights will be available.
@@ 2 Related Work
Multi-Stage WSSS. The majority of image-level WSSS work is multi-stage, typically comprising three stages: classification training (CAM generation), CAM refinement, and segmentation training. Some approaches employ heuristic strategies to address incomplete activation regions. For example, adversarial erasing [71, 58, 30, 68], feature map optimization [32, 14, 13, 12], self-supervised learning [60, 11, 56], and contrastive learning [27, 64, 73, 15] are employed. Some methods focus on post-refining the CAMs by propagating object regions from the seeds to their semantically similar pixels. AffinityNet [2], for instance, learns pixel-level affinity to enhance CPL. This has motivated other work [1, 20, 9, 38] that utilize additional networks to generate more accurate CPL. Other work is dedicated to studying optimization given the coarse pseudo-labels: [40] explores uncertainty of noisy labels, [43] adaptively corrects CPL during early learning, and [50] enhances boundary prediction through co-training. Since image-level labels alone do not yield satisfactory results, several methods incorporate additional modalities, such as saliency maps [37, 38, 73, 18] and CLIP models [67, 42, 63]. More recently, vision transformers [17] have emerged as prominent models for various vision tasks. Several WSSS studies benefit from vision transformers: [21] enhance CAMs by incorporating the attention map from ViT; [65] introduces class-specific attention for discriminative object localization; [42] and [67] leverage multi-modal transformers to enhance performance.
Single-Stage WSSS. In contrast, single-stage methods are much more efficient. They contain a shared backbone with heads for classification and segmentation [3, 70, 52, 53]. The pipeline involves generating and refining the CAMs, leveraging an offline module, such as PAMR [3], PAR [52], or CRF [70]. Subsequently, the refined CPL are used for segmentation. Single-stage methods exhibit faster speed and a lower memory footprint but are challenging to optimize due to the obfuscation in offline refinement. As a result, they often yield inferior performance compared to multi-stage methods. More recently, with the success of ViT, single-stage WSSS has been greatly advanced. AFA [52] proposes learning reliable affinity from attention to refine the CAMs. Similarly, ToCo [53] mitigates the problem of over-smoothing in vision transformers by contrastively learning from patch tokens and class tokens.
The existing works depend heavily on offline refinement of CAMs. In this study, we further explore the potential of single-stage approaches and showcase the redundancy of offline refinement. We propose an effective alternative for generating consistent, and accurate CAMs in WSSS.
@@ 3.1 Guiding Class Activation Maps
Figure 1: Oracle Experiments on VOC. CAMs are guided by the ground truth (GT), proposed segmentation pseudo-labels (SPL), no guidance (NO) and random noise (NS). (a): classification performance; (b): CAM quality; (c) CAMs visualization. All experiments employ 2k-iters warm-up before guidance is introduced.
To demonstrate this, we conduct oracle experiments in which we supervise the output CAMs from a classifier directly with the ground truth segmentation (GT). This enables all elements in $F$ to be optimized. For comparison, we conduct experiments wherein the CAMs are (i) not guided (NO), (ii) guided with masks of random noise (NS).
Results, shown in Fig. 1, demonstrate that different guidance for $M$ does not affect classification even for the NS group, as all experiment groups achieved over 97% classification precision. However, drastic differences can be observed w.r.t. the quality of the CAMs. The GT group results in a notable quality improvement over the NO group, as shown in Fig. 1(b)(c). In contrast, the NS group sabotages the CAMs. This suggests the stochasticity of CAMs and explains their variability and lack robustness, something also observed in [2, 60, 13].
@@
Since relying on GT segmentation is not feasible in WSSS, we propose an alternative for guiding CAMs, employing mask predictions as segmentation pseudo-labels (SPL). As shown in Fig. 1, an SPL-guided classifier yields CAMs that significantly outperform vanilla CAMs (NO group), performing close to the oracle trained with the GT. Motivated by this, we introduce a co-training mechanism in which CAMs and mask predictions are optimized mutually without any additional CAM refinement.
3.2 Co-training with Swapping Assignments
@@
Swapping Assignments. Our objective is to co-optimize $\mathcal{S}$ and $\mathcal{M}$. A naive approach could enforce the learning objectives $\mathcal{S}\triangleq\mathcal{S}^{\prime}$ and $\mathcal{M}\triangleq\mathcal{M}^{\prime}$ as a knowledge distillation process [25], where AN and ON play the roles of teacher and student. However, this assumes availability of a pretrained teacher which is not possible in WSSS settings. Instead, we setup a swapped self-distillation with the objective:
$\small\mathcal{L}_{\text{swap}}=\mathcal{L}_{\text{c2s}}(\mathcal{S},\mathcal{M}^{\prime})+\mathcal{L}_{\text{s2c}}(\mathcal{M},\mathcal{S}^{\prime})~{},\vspace{-5pt}$ (2)
where $\mathcal{L}_{\text{c2s}}$ optimizes the segmentation performance given the CPL, and $\mathcal{L}_{\text{s2c}}$ considers the CAM quality with respect to SPL. Building on self-distillation [6, 48], we present this swapped self-distillation framework tailored to facilitate information exchange between the CAMs and segmentation.
Figure 3: CPL Analysis on val splits of VOC (a, c, e) and COCO (b, d, f). (a) and (b): heatmap of CPL accuracy vs. confident ranges (x-axis) for different time-steps (y-axis). (c) and (d): correlation between perplexity and accuracy of CPL for different time-steps. (e) and (f): distribution of CAMs’ confidence categorized by the proposed dynamic threshold. (best viewed under zoom)
3.3 Segmentation Optimization.
CAM2Seg Learning. As the CAMs in CoSA are inherently guided, extra refinement [29, 2, 52] is not required, and they can be directly employed as learning targets. Nonetheless, CAMs primarily concentrate on the activated regions of the foreground while disregarding the background. As per the established convention [60, 15, 53], a threshold value $\xi$ is employed for splitting the foreground and the background. Formally, the CAM pseudo-label (CPL) is given by:
$\small\hat{\mathcal{Y}}^{\text{CPL}}_{x,y}=\left\{\begin{aligned} &\mathtt{argmax}(\mathcal{M}_{x,y}^{\prime})+1,&\text{if $\nu\geq\xi$,}\\ &0,&\text{if $\nu<\xi$,}\\ \end{aligned}\right.,\vspace{-5pt}$ (3)

### v2

@@ Abstract
Class activation maps (CAMs) are commonly employed in weakly supervised semantic segmentation (WSSS) to produce pseudo-labels. Due to incomplete or excessive class activation, existing studies often resort to offline CAM refinement, introducing additional stages or proposing offline modules. This can cause optimization difficulties for single-stage methods and limit generalizability. In this study, we aim to reduce the observed CAM inconsistency and error to mitigate reliance on refinement processes. We propose an end-to-end WSSS model incorporating guided CAMs, wherein our segmentation model is trained while concurrently optimizing CAMs online. Our method, Co-training with Swapping Assignments (CoSA), leverages a dual-stream framework, where one sub-network learns from the swapped assignments generated by the other. We introduce three techniques in this framework: i) soft perplexity-based regularization to penalize uncertain regions; ii) a threshold-searching approach to dynamically revise the confidence threshold; and iii) contrastive separation to address the coexistence problem. CoSA demonstrates exceptional performance, achieving mIoU of 76.2% and 51.0% on VOC and COCO validation datasets, respectively, surpassing existing baselines by a substantial margin. Notably, CoSA is the first single-stage approach to outperform all existing multi-stage methods including those with additional supervision. Source code is publicly available at here.
Keywords:
Weakly-supervised Learning Semantic Segmentation CAM
@@ 1 Introduction
The objective of weakly supervised semantic segmentation (WSSS) is to train a segmentation model without relying on pixel-level labels but on weak and cost-effective annotations, such as image-level classification labels [3, 26, 50], object points [4, 45], and bounding boxes [16, 28, 57, 36]. In particular, image-level classification labels have commonly been employed as weak labels due to the minimal or negligible annotation effort required [2, 60]. With the absence of precise localization information, image-level WSSS often makes use of the coarse localization offered by class activation maps (CAMs) [75]. CAMs pertain to the intermediate outputs derived from a classification network. They visually illustrate the activation regions corresponding to each individual class. Thus, they are often used to generate pseudo masks for training. However, CAMs suffer from i) Inconsistent Activation: CAMs demonstrate variability and lack robustness in accommodating geometric transformations of input images [60], resulting in inconsistent activation regions for the same input. ii) Inaccurate Activation: activation region accuracy is often compromised, resulting in incomplete or excessive class activation, only covering discriminative object regions [1]. Despite enhanced localization mechanisms in the variants GradCAM [55] and GradCAM++ [7], they still struggle to generate satisfactory pseudo-labels for WSSS [60]. Thus, many WSSS works are dedicated to studying CAM refinement or post-processing [1, 15, 31].
@@
In general, they [2, 65, 18, 50] comprise three stages: CAM generation, refinement, and segmentation training with pseudo-labels. Multi-stage frameworks are known to be time-consuming and complex as several models must be trained at different stages. In contrast, single-stage models [3, 73, 52], which include a unified network of all stages, are more efficient. They are trained to co-optimize the segmentation and classification tasks, but the generated CAMs are not explicitly trained. As a result, they need refinement to produce high-quality pseudo-labels, often leveraging hand-crafted modules, such as CRF in [73], PAMR in [3], PAR in [52, 53]. As the refinement modules are predefined and offline, they decouple the CAMs from the primary optimization. When the refined CAMs are employed as segmentation learning objectives, the optimization of the segmentation branch may deviate from that of the classification branch. Hence, it is difficult for single-stage models to optimize the segmentation task while yielding satisfactory CAM pseudo-labels. This optimization difficulty underlies the inferior performance in single-stage approaches compared to multi-stage [65, 50]. Further, hand-crafted refinement modules require heuristic tuning and empirical changes, thereby limiting their adaptability to novel datasets [3, 52]. Despite the potential benefits of post-refinement in addressing the aforementioned issues associated with CAMs, which have been extensively discussed in WSSS studies, there has been limited exploration of explicit online optimization for CAMs.
@@
The absence of fully optimized CAMs is an important factor in the indispensability of this refinement. In this paper, we take a different approach by optimizing CAMs in an end-to-end fashion. We ask a core question: Can we train a model that delivers reliable, consistent and accurate CAMs, which can be applied directly for WSSS without the necessity for subsequent refinements? We show that the answer is yes, in two respects: 1) we note that even though CAM is differentiable, it is not robust to variation. As the intermediate output of classification, CAMs are not fully optimized for segmentation purposes since the primary objective is to minimize classification error. This implies that within an optimized network, numerous weight combinations exist that can yield accurate classification outcomes, while generating CAMs of varying qualities. To investigate this, we conduct oracle experiments, training a classification model while simultaneously guiding the CAMs with the segmentation ground truth. A noticeable enhancement in quality is observed in guided compared to vanilla CAMs, without compromising classification accuracy. 2) we demonstrate the feasibility of substituting the oracle with segmentation pseudo-labels (SPL) in the context of weak supervision. Consequently, we harness the potential of SPL for WSSS by co-training both CAMs and segmentation through mutual learning.
@@
We explore an effective way to substitute the CAM refinement process, i.e. guiding CAMs in an end-to-end fashion. Our method optimizes the CAMs and segmentation prediction simultaneously thanks to the differentiability of CAMs. To achieve this, we adopt a dual-stream framework that includes an online network (ON) and an assignment network (AN), inspired by self-training frameworks [5, 22, 68]. The AN is responsible for producing CAM pseudo-labels (CPL) and segmentation pseudo-labels (SPL) to train the ON. Since CPL and SPL are swapped for supervising segmentation and CAMs, respectively, our method is named Co-training with Swapping Assignments (CoSA).
The benefit of this end-to-end framework is that it enables us to quantify pseudo-label reliability online, as opposed to the offline hard pseudo-labels used in existing methodologies [2, 65, 15, 50]. We can then incorporate soft regularization to compensate for CPL uncertainty, where the segmentation loss for different regions is adaptively weighted according to our estimated perplexity map. In comparison to existing literature, this dynamic learning scheme can exploit the potential of CPL and enhance the final performance, as opposed to performance being constrained by predetermined CPL. The threshold is a key hyper-parameter for generating the CPL [60, 50, 53]. It not only requires tuning but necessitates dynamic adjustment to align with the model’s learning state at various time-steps. CoSA integrates threshold searching to dynamically adapt its learning state, as opposed to the fixed thresholding [18, 12, 52]. This can enhance performance and help to eliminate the laborious manual parameter-tuning task. We further address a common issue with CAMs, known as the coexistence problem, whereby certain class activations display extensive false positive regions that inaccurately merge the objects with their surroundings (Fig. 4). In response, we introduce a technique to leverage low-level CAMs enriched with object-specific details to contrastively separate those coexistent classes.
@@
The proposed CoSA greatly surpasses existing WSSS methods. Our approach achieves the leading results on VOC and COCO benchmarks, highlighting the contribution of this work: i) We are the first to propose SPL as a substitute for guiding CAMs. We present compelling evidence of its potential to produce more reliable, consistent and accurate CAMs. ii) We introduce a dual-stream framework with swapped assignments, which co-optimizes the CAMs and segmentation predictions in an end-to-end fashion. iii) We address the learning dynamics, proposing two components within our framework: reliability-based adaptive weighting and dynamic thresholding. iv) We address the CAM coexistence issue, proposing a contrastive separation approach to regularize CAMs, significantly enhancing the results of affected classes.
@@ 2 Related Work
Multi-Stage WSSS.Most image-level WSSS work is multi-stage, typically comprising three stages: CAM generation, CAM refinement, and segmentation training. Some approaches employ heuristic strategies to address incomplete activation regions, such as adversarial erasing [74, 58, 30, 70], feature map optimization [32, 14, 12, 13], self-supervised learning [60, 11, 56], and contrastive learning [27, 64, 76, 15]. Some methods focus on post-refining the CAMs by propagating object regions from the seeds to their semantically similar pixels. AffinityNet [2], for instance, learns pixel-level affinity to enhance CPL. This has motivated other work [1, 20, 10, 38] that utilize additional networks to generate more accurate CPL. Other work studies optimization given coarse pseudo-labels: [40] explores uncertainty of noisy labels, [43] adaptively corrects CPL during early learning, and [50] enhances boundary prediction through co-training. Since image-level labels alone do not yield satisfactory results, several methods incorporate additional modalities, such as saliency maps [37, 38, 76, 18] or CLIP models [67, 42, 63]. Recently, vision transformers [17] have emerged as prominent models for various vision tasks. Several WSSS studies benefit from vision transformers: [21] enhances CAMs by incorporating the attention map from ViT; [65] introduces class-specific attention for discriminative object localization; [42] and [67] leverage multi-modal transformers to enhance performance.
Single-Stage WSSS. In contrast, single-stage methods are much more efficient. They contain a shared backbone with heads for classification and segmentation [3, 73, 52, 53]. The pipeline involves generating and refining the CAMs, leveraging an offline module, such as PAMR [3], PAR [52], or CRF [73]. Subsequently, the refined CPL are used for segmentation. Single-stage methods exhibit faster speed and a lower memory footprint but are challenging to optimize due to the obfuscation in offline refinement. As a result, they often yield inferior performance compared to multi-stage methods. More recently, with the success of ViT, single-stage WSSS has been greatly advanced. AFA [52] proposes learning reliable affinity from attention to refine the CAMs. Similarly, ToCo [53] mitigates the problem of over-smoothing in vision transformers by contrastively learning from patch tokens and class tokens. The existing works depend heavily on offline refinement of CAMs. In this study, we further explore the potential of single-stage approaches and showcase the redundancy of offline refinement. We propose an effective alternative for generating consistent, and accurate CAMs in WSSS.
@@ 3.1 Guiding Class Activation Maps
To demonstrate, we conduct oracle experiments wherein we supervise the output CAMs from a classifier with the ground truth segmentation (GT), enabling optimization of all elements in $F$. For comparison, we conduct experiments where the CAMs are not guided (NO), and guided with random noise masks (NS). Results, shown in Fig. 1, demonstrate that different guidance for $M$ does not affect classification even for the NS group, as all experiment groups achieved over 97% classification precision. However, drastic differences can be observed w.r.t. the quality of the CAMs. The GT group results in a notable quality improvement over the NO group, as shown in Fig. 1(b)(c). In contrast, the NS group sabotages the CAMs. This suggests the stochasticity of CAMs and explains their variability and lack robustness, something also observed in [2, 60, 12].
Since relying on GT segmentation is not feasible in WSSS, we propose an alternative for guiding CAMs, employing predicted masks as segmentation pseudo-labels (SPL). As shown in Fig. 1, a SPL-guided classifier yields CAMs that significantly outperform vanilla CAMs (NO), performing close to the oracle (GT). Motivated by this, we introduce a co-training mechanism in which CAMs and predicted masks are optimized mutually without additional CAM refinement.
Figure 1: Oracle Experiments on VOC. CAMs are guided by the ground truth (GT), proposed segmentation pseudo-labels (SPL), no guidance (NO) and random noise (NS). (a): classification performance; (b): CAM quality; (c) CAM visualization. All experiments employ 2k-iters warm-up before guidance is introduced.
3.2 Co-training with Swapping Assignments
@@
Swapping Assignments. Our objective is to co-optimize $\mathcal{S}$ and $\mathcal{M}$. A naive approach could enforce the learning objectives $\mathcal{S}\triangleq\mathcal{S}^{\prime}$ and $\mathcal{M}\triangleq\mathcal{M}^{\prime}$ as a knowledge distillation process [25], where AN and ON play the roles of teacher and student. However, this assumes availability of a pretrained teacher which is not possible in WSSS settings. Instead, we setup a swapped self-distillation objective:
$\small\mathcal{L}_{\text{swap}}=\mathcal{L}_{\text{c2s}}(\mathcal{S},\mathcal{M}^{\prime})+\mathcal{L}_{\text{s2c}}(\mathcal{M},\mathcal{S}^{\prime})~{},$ (2)
where $\mathcal{L}_{\text{c2s}}$ optimizes the segmentation performance given the CPL, and $\mathcal{L}_{\text{s2c}}$ assesses the CAM quality with respect to SPL. Building on self-distillation [6, 48], we present this swapped self-distillation framework tailored specifically to facilitate information exchange between the CAMs and segmentation.
3.3 Segmentation Optimization
CAM2Seg Learning. Previous studies [29, 2, 3, 52] refine the CAMs to obtain pseudo-label, then perform pseudo-label to segmentation learning (PL2Seg). As our guided-CAMs do not require extra refinement process, they can be directly employed as learning targets (CAM2Seg). Nonetheless, CAMs primarily concentrate on the activated regions of the foreground while disregarding the background. As per the established convention [60, 15, 53], a threshold value $\xi$ is employed for splitting the foreground and the background. Formally, our CAM pseudo-label (CPL) is given by:
$\small\hat{\mathcal{Y}}^{\text{CPL}}_{x,y}=\left\{\begin{aligned} &\mathtt{argmax}(\mathcal{M}_{x,y}^{\prime})+1,&\text{if $\nu\geq\xi$,}\\ &0,&\text{if $\nu<\xi$,}\\ \end{aligned}\right.,$ (3)